Dadayeva Inc. has $5 million of 6% convertible bonds outstanding. Each $1,000 bond is convertible into 50 no par value common shares. The bonds pay interest on January 31 and July 31. On July 31, 2020, the holders of $1,250,000 of these bonds exercised the conversion privilege. On that date, the market price of the bonds was 110, the market price of the common shares was $40, the carrying value of the common shares was $20, and the Contributed Surplus—Conversion Rights account balance was $700,000. The total unamortized bond premium at the date of conversion was $362,990. The remaining bonds were never converted and were retired when they reached the maturity date. Assume that the company follows IFRS.

Instructions

a.  

Assuming that the book value method was used, record the conversion of the $1,250,000 of bonds on July 31, 2020.

b.  

Prepare the journal entry that would be required for the remaining amount in Contributed Surplus—Conversion Rights when the maturity of the remaining bonds is recorded.
Answer #1

a)

The book value of convertible bonds converted =  $1250000 +[ $362990* $1250000 / 5000000 ] - $1250000/1000*50*20 = 1250000+90747.50-1250000= $90747.50

b)

Bonds payable Dr $1250000
Premium on bonds payable $90747.50
To Common stock $1250000
To Paid up capital in excess of par $90747.50